# Free Descriptions of Convex Sets

### v2

@@ 4.3 Permutahedra, Schur-Horn orbitopes, and Limits
We remark that the above constructions can be extended to handle vectors and matrices of all sizes, not just of sizes $(m2^{n})_{n\in\mathbb{N}}$. This can be done by indexing our consistent sequences by posets. In particular, our results generalize to the following, more complicated, sequences of group representations. If $\mathcal{N}$ is a strict poset, an $\mathcal{N}$-indexed consistent sequence of $\{\mathsf{G}_{n}\}_{n\in\mathcal{N}}$-representations is a sequence $\{(\mathbb{\mathbb{missing}}V_{n},\varphi_{N,n})\}_{n<N\in\mathcal{N}}$ of $\mathsf{G}_{n}$-representations and embeddings $\varphi_{N,n}\colon\mathbb{\mathbb{missing}}V_{n}\hookrightarrow\mathbb{\mathbb{missing}}V_{N}$ for each $n<N$ such that $\varphi_{N,n}$ is $\mathsf{G}_{n}$-equivariant, and $\varphi_{M,N}\circ\varphi_{N,n}=\varphi_{M,n}$ whenever $n<N<M$. All our results apply in this setting after replacing all occurrences of $n+1$ by $N>n$. To handle permutahedra and Schur-Horn orbitopes of any sizes we let $\mathcal{N}=\mathbb{N}$ with the divisibility partial order, whereby $n\leq m$ iff $m=nk$ for some $k\in\mathbb{N}$, with embeddings $\varphi_{kn,n}\colon\mathbb{R}^{n}\hookrightarrow\mathbb{R}^{nk}$ sending $x\mapsto x\otimes\mathbbm{1}_{k}$ or $\varphi_{kn,n}\colon\mathbb{S}^{n}\hookrightarrow\mathbb{S}^{nk}$ sending $X\mapsto X\otimes I_{k}$.
@@ 4.4 Free Spectrahedra
As the only morphisms $\mathscr{V}_{0}\to\mathscr{V}_{0}$ are multiples of the identity, and elements of $\mathbb{S}^{k}\otimes\mathbb{S}^{n}$ can be viewed as $k\times k$ symmetric block matrices with symmetric $n\times n$ blocks of $n\times n$ symmetric matrices, we conclude that the morphisms $\mathscr{V}\to\mathscr{U}$ are precisely maps of the form $(X_{1},\ldots,X_{d})\mapsto\sum_{i}L_{i}\otimes X_{i}$ for some $L_{1},\ldots,L_{d}\in\mathbb{S}^{k}$. Note that sequences of adjoints of such maps are also morphisms in this case. As the only $\mathsf{G}_{n}$-invariants in $\mathbb{S}^{n}$ are multiples of $I_{n}$, the space of freely-described elements in $\mathscr{U}$ is $\{\{L_{0}\otimes I_{n}\}_{n}:L_{0}\in\mathbb{S}^{k}\}$, which satisfy the condition $L_{0}\otimes(I_{n+1}-I_{n})\succeq 0$ from Proposition 3.2(a) if and only if $L_{0}\succeq 0$. Thus, the parametric family of free descriptions certifying compatibility as in Proposition 3.2(a) is
@@
By [72, §2.3], it suffices to show that if $(X_{1},\ldots,X_{d})\in\mathcal{C}_{n}$ and $(Y_{1},\ldots,Y_{d})\in\mathcal{C}_{m}$ then $(X_{1}\oplus Y_{1},\ldots,X_{d}\oplus Y_{d})\in\mathcal{C}_{n+m}$, and that $(V^{\top}X_{1}V,\ldots,V^{\top}X_{d}V)\in\mathcal{C}_{k}$ for any isometry $V\in\mathbb{R}^{n\times k}$. For the first, intersection compatibility shows that $(X_{1}\oplus 0_{m},\ldots,X_{d}\oplus 0_{m})\in\mathcal{C}_{n+m}$ and $(Y_{1}\oplus 0_{n},\ldots,Y_{d}\oplus 0_{n})\in\mathcal{C}_{n+m}$. Conjugating the latter tuple by appropriate permutation matrices, we conclude that $(0_{n}\oplus Y_{1},\ldots,0_{n}\oplus Y_{d})\in\mathcal{C}_{n+m}$. Finally, since $\mathcal{C}_{n+m}$ is a convex cone we get $(X_{1}\oplus Y_{1},\ldots,X_{d}\oplus Y_{d})\in\mathcal{C}_{n+m}$, so $\{\mathcal{C}_{n}\}$ is closed under direct sums. For the second, any isometry $V\in\mathbb{R}^{n\times k}$ can be written as $V=U\iota_{n,k}$ where $U\in\mathsf{O}_{n}$ and $\iota_{n,k}\colon\mathbb{R}^{k}\to\mathbb{R}^{n}$ is the zero-padding embedding. Observe that $(\iota_{n,k}^{\top}X_{1}\iota_{n,k},\ldots,\iota_{n,k}^{\top}X_{d}\iota_{n,k})=\mathcal{P}_{k}(X_{1},\ldots,X_{d})$ is the orthogonal projection of $(X_{1},\ldots,X_{d})$ onto $\mathbb{V}_{k}$ when embedded in $\mathbb{V}_{n}$. Thus, $(V^{\top}X_{1}V,\ldots,V^{\top}X_{d}V)=\mathcal{P}_{k}(U^{\top}X_{1}U,\ldots,U^{\top}X_{d}U)\in\mathcal{C}_{k}$ since $\mathcal{C}_{n}$ is $\mathsf{O}_{n}$-invariant and since $\{\mathcal{C}_{n}\}$ is projection-compatible. ∎
@@ 4.5 Spectral Functions, (Quantum) Entropy, and Variants
Here $\mathrm{dom}(h_{n})=\Delta^{n-1}$ and $\mathrm{dom}(D_{n})=(\mathbb{R}^{n}_{+})^{2}$, while $\mathrm{dom}(\mathsf{H}_{n})=\mathcal{D}^{n-1}$ is the spectraplex from Example 1.1(b) and $\mathrm{dom}(S_{n})=(\mathbb{S}^{n}_{+})^{2}$. We use the standard convention that $00\log\frac{0}{y}=$ 0 even if $y=0$, and $x\log\frac{x}{0}=\infty$ when $x\neq 0$ [74, §2.3]. These sequences of functions are intersection-compatible but not projection-compatible (e.g., their domains are not projection-compatible).
@@
$E_{n}^{(m,k)}$ =$=\Bigg{\{}(X,t)\in\mathbb{S}^{n}\oplus\mathbb{R}\Bigg{|}\exists T_{0},\ldots,T_{m},Z_{0},\ldots,Z_{k}\in\mathbb{S}^{n}\textrm{ s.t. }Z_{0}=I_{n},\ \sum_{j=1}^{m}w_{j}T_{j}=-2^{-k}T_{0},$
$\quad\begin{bmatrix}Z_{i}&Z_{i+1}\\ Z_{i+1}&X\end{bmatrix}\succeq 0,\textrm{ for }i=0,\ldots,k-1,\ \begin{bmatrix}Z_{k}-X-T_{j}&-\sqrt{s_{j}}T_{j}\\ -\sqrt{s_{j}}T_{j}&X-s_{j}T_{j}\end{bmatrix}\succeq 0,$
$\quad\textrm{for }j=1,\ldots,m,\ \mathrm{Tr}(T_{0})\leq t\Bigg{\}},$
@@
This is a free description of the form (ConicSeq) which almost, but not quite, satisfies the conditions of Proposition 3.2. Indeed, let $\mathscr{W}=\mathscr{V}^{\oplus(m+k+1)}$ and $\mathscr{U}=\mathscr{V}\oplus(\mathbb{S}^{2}\otimes\mathscr{V})^{\oplus(m+k)}\oplus\mathbb{R}$ containing the cones $\{\mathcal{K}_{n}=\{0\}\oplus(\mathbb{S}^{2}\otimes\mathbb{S}^{n})_{+}^{\oplus(m+k)}\oplus\mathbb{R}_{+}\}$. Define
@@
$\qquad\bigoplus_{j=1}^{m}\left(\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes Z_{k}-\begin{bmatrix}1&\sqrt{s_{j}}\\ \sqrt{s_{j}}&s_{j}\end{bmatrix}\otimes T_{j}\right),-\mathrm{Tr}(T_{0})\Bigg{)},$
@@
Note that $\{A_{n}\}$ and $\{B_{n}\}$ are morphisms, that $\{u_{n}\}$ is a freely-described element of $\mathscr{U}$ satisfying $u_{n+1}-u_{n}\in\mathcal{K}_{n+1}$, but that $\{B_{n}^{*}\}$ is not a morphism because of the $\mathrm{Tr}(T_{0})$ term in $B_{n}$. By the proof of Proposition 3.2, this description certifies that $E_{n}^{(m,k)}\subseteq E_{n+1}^{(m,k)}$ but not $E_{n+1}^{(m,k)}\cap(\mathbb{S}^{n}\oplus\mathbb{R})\subseteq E_{n}^{(m,k)}$. Analogously to the naïve description of the solid simplex in Section 4.1, this free description does not make the intersection compatibility of $\{E_{n}^{(m,k)}\}$ obvious.
@@
Note that $\mathbb{S}^{2}\otimes\mathbb{S}^{n}\cong(\mathbb{S}^{n})^{3}$ as $\mathsf{O}_{n}$-representations, and these isomorphisms commute with zero-padding, so that $\mathscr{U}\cong\mathscr{V}^{1+3(m+k)}\oplus\mathbb{R}$ as consistent sequences. As $\dim(\mathbb{S}^{n})^{\mathsf{O}_{n}}=1$ and $\dim\mathcal{L}(\mathbb{S}^{n})^{\mathsf{O}_{n}}=2$, the dimension of invariants parametrizing free descriptions are
$\dim\mathcal{L}(\mathbb{V}_{n},\mathbb{U}_{n})^{\mathsf{O}_{n}}=3[2(m+k)+1],\quad\dim\mathcal{L}(\mathbb{W}_{n},\mathbb{U}_{n})^{\mathsf{O}_{n}}=3(m+k+1)[2(m+k)+1],$
@@ 4.6 Graph Parameters
Let $\mathscr{V}=\{\mathbb{S}^{n}\}$ with embeddings by zero-padding and the action of $\mathsf{G}_{n}=\mathsf{S}_{n}$ by conjugation. As discussed in Section 1.1.2 and Example 1.7, graph parameters are sequences $\{f_{n}\colon\mathbb{S}^{n}\to\mathbb{R}\}$ of $\mathsf{G}_{n}$-invariant functions, and many standard graph parameters are convex (or concave) and satisfy either intersection or projection compatibility. Furthermore, it is desirable to find relaxations for standard graph parameters satisfying the same compatibility conditions. We consider here the examples of max-cut and inverse stability number, and we show that their usual relaxations are freely-described with descriptions certifying their compatibility. We then derive parametric families of free descriptions from our framework that may be used to fit graph parameters to data.
@@
The sequence $\{\mathcal{E}_{n}\}$ also satisfies projection compatibility, and the above is a free description certifying this compatibility. Indeed, let $\mathscr{W}=\{0\}$ and $\mathscr{U}=\mathscr{V}\oplus\{\mathbb{R}^{n}\}$ where the latter sequence is the usual one from Example 1.5, with cones $\mathscr{K}=\{\mathbb{S}^{n}_{+}\oplus\{0\}\}$. Then (17) is of the form (ConicSeq) with $A_{n}X=(X,\mathrm{diag}(X))$, $B_{n}=0$, and $u_{n}=(0,-\mathbbm{1}_{n})$. Note that $\{A_{n}\},\{A_{n}^{*}\}$ are morphisms and $u_{n+1}-u_{n}=(0,-e_{n+1})=(e_{n+1}e_{n+1}^{\top},0)-A_{n+1}e_{n+1}e_{n+1}^{\top}$, hence Proposition 3.2(b) applies. Neither the cut polytopes nor the elliptopes is intersection-compatible, as zero-padding a matrix with all-1’s diagonal does not yield such a matrix. The sequences $\{\mathrm{CUT}_{n}-I_{n}\}$ and $\{\mathcal{E}_{n}-I_{n}\}$ are, however, both intersection- and projection-compatible, and the shifted elliptopes admit free descriptions certifying their compatibility.
@@
where $X\geq 0$ denotes an entrywise nonnegative matrix. Both $\{\mathcal{D}_{n}\}$ and $\{\widetilde{\mathcal{D}}_{n}\}$ are intersection-compatible. Moreover, the above description of $\{\widetilde{\mathcal{D}}_{n}\}$ is free and certifies this compatibility as in Proposition 3.2(b). Indeed, let $\mathscr{W}=\{0\}$ and $\mathscr{U}=\mathscr{V}\oplus\mathbb{R}$ with cones $\mathcal{K}_{n}=(\mathbb{S}^{n}_{+}\cap\mathbb{R}^{n\times n}_{+})\oplus 0$. Then the above description of $\widetilde{\mathcal{D}}_{n}$ is of the form (ConicSeq) with $A_{n}X=(X,\mathbbm{1}_{n}^{\top}X\mathbbm{1}_{n})$, $B_{n}=0$, and $u_{n}=(0,-1)$. Note that $\{A_{n}\}$ is a morphism and $u_{n}=u_{n+1}$, hence Proposition 3.2(a) applies. Neither $\mathcal{D}_{n}$ nor $\widetilde{\mathcal{D}}_{n}$ is projection-compatible since their support functions, which are the inverse stability number and its semidefinite approximation above, are not intersection compatible (see Section 1.3). Indeed, appending isolated vertices to a graph increases its stability number.
@@
Beyond the two particular preceding examples, we derive here an expressive parametric family of convex graph parameters; in addition to above examples, our family also includes several other examples from [76]. To that end, let $\mathscr{W}=\mathscr{U}=\mathrm{Sym}^{2}(\mathrm{Sym}^{\leq 2}\{\mathbb{R}^{n}\})=\{\mathbb{S}^{\binom{n+2}{2}}\}$. The dimensions of invariants parametrizing free descriptions in this case is too large for us to write explicit bases for them as functions of $n$. Instead, we compute these dimensions using the algorithm in Section 6.1 below, see Example 6.2(b):
@@
$\dim\Big{\{}\{A_{n}\colon\mathbb{V}_{n}\to\mathbb{U}_{n}\}\textrm{ morphism}\Big{\}}=19,$ (18)
$\dim\Big{\{}\{B_{n}\colon\mathbb{W}_{n}\to\mathbb{U}_{n}\}:\textrm{both }\{B_{n}\}\textrm{ and }\{B_{n}^{*}\}\textrm{ morphisms}\Big{\}}=104.$

### v3

@@ 4.3 Permutahedra, Schur–Horn Orbitopes, and Limits
We remark that the above constructions can be extended to handle vectors and matrices of all sizes, not just of sizes $(m2^{n})_{n\in\mathbb{N}}$. This can be done by indexing our consistent sequences by posets. In particular, our results generalize to the following, more complicated, sequences of group representations. If $\mathcal{N}$ is a strict poset, an $\mathcal{N}$-indexed consistent sequence of $\{\mathsf{G}_{n}\}_{n\in\mathcal{N}}$-representations is a sequence $\{(\mathbb{\mathbb{}}V_{n},\varphi_{N,n})\}_{n<N\in\mathcal{N}}$ of $\mathsf{G}_{n}$-representations and embeddings $\varphi_{N,n}\colon\mathbb{\mathbb{}}V_{n}\hookrightarrow\mathbb{\mathbb{}}V_{N}$ for each $n<N$ such that $\varphi_{N,n}$ is $\mathsf{G}_{n}$-equivariant, and $\varphi_{M,N}\circ\varphi_{N,n}=\varphi_{M,n}$ whenever $n<N<M$. All our results apply in this setting after replacing all occurrences of $n+1$ by $N>n$. To handle permutahedra and Schur–Horn orbitopes of any sizes we let $\mathcal{N}=\mathbb{N}$ with the divisibility partial order, whereby $n\leq m$ iff $m=nk$ for some $k\in\mathbb{N}$, with embeddings $\varphi_{kn,n}\colon\mathbb{R}^{n}\hookrightarrow\mathbb{R}^{nk}$ sending $x\mapsto x\otimes\mathbbm{1}_{k}$ or $\varphi_{kn,n}\colon\mathbb{S}^{n}\hookrightarrow\mathbb{S}^{nk}$ sending $X\mapsto X\otimes I_{k}$.
@@ 4.4 Free Spectrahedra
As the only morphisms $\mathscr{V}_{0}\to\mathscr{V}_{0}$ are multiples of the identity, and elements of $\mathbb{S}^{k}\otimes\mathbb{S}^{n}$ can be viewed as $k\times k$ symmetric block matrices with symmetric $n\times n$ blocks of $n\times n$ symmetric matrices, we conclude that the morphisms $\mathscr{V}\to\mathscr{U}$ are precisely maps of the form $(X_{1},\ldots,X_{d})\mapsto\sum_{i}L_{i}\otimes X_{i}$ for some $L_{1},\ldots,L_{d}\in\mathbb{S}^{k}$. Note that sequences of adjoints of such maps are also morphisms in this case. As the only $\mathsf{G}_{n}$-invariants in $\mathbb{S}^{n}$ are multiples of $I_{n}$, the space of freely-described elements in $\mathscr{U}$ is $\{\{L_{0}\otimes I_{n}\}_{n}:L_{0}\in\mathbb{S}^{k}\}$, which satisfy the condition $L_{0}\otimes(I_{n+1}-I_{n})\succeq 0$ from Proposition 3.2(a) if and only if $L_{0}\succeq 0$. Thus, the parametric family of dimension-free descriptions certifying compatibility as in Proposition 3.2(a) is
@@
By [70, §2.3], it suffices to show that if $(X_{1},\ldots,X_{d})\in\mathcal{C}_{n}$ and $(Y_{1},\ldots,Y_{d})\in\mathcal{C}_{m}$ then $(X_{1}\oplus Y_{1},\ldots,X_{d}\oplus Y_{d})\in\mathcal{C}_{n+m}$, and that $(V^{\top}X_{1}V,\ldots,V^{\top}X_{d}V)\in\mathcal{C}_{k}$ for any isometry $V\in\mathbb{R}^{n\times k}$. For the first, intersection compatibility shows that $(X_{1}\oplus 0_{m},\ldots,X_{d}\oplus 0_{m})\in\mathcal{C}_{n+m}$ and $(Y_{1}\oplus 0_{n},\ldots,Y_{d}\oplus 0_{n})\in\mathcal{C}_{n+m}$. Conjugating the latter tuple by appropriate permutation matrices, we conclude that $(0_{n}\oplus Y_{1},\ldots,0_{n}\oplus Y_{d})\in\mathcal{C}_{n+m}$. Finally, since $\mathcal{C}_{n+m}$ is a convex cone we get $(X_{1}\oplus Y_{1},\ldots,X_{d}\oplus Y_{d})\in\mathcal{C}_{n+m}$, so $\{\mathcal{C}_{n}\}$ is closed under direct sums. For the second, any isometry $V\in\mathbb{R}^{n\times k}$ can be written as $V=U\iota_{n,k}$ where $U\in\mathsf{O}_{n}$ and $\iota_{n,k}\colon\mathbb{R}^{k}\to\mathbb{R}^{n}$ is the zero-padding embedding. Observe that $(\iota_{n,k}^{\top}X_{1}\iota_{n,k},\ldots,\iota_{n,k}^{\top}X_{d}\iota_{n,k})=\mathcal{P}_{k}(X_{1},\ldots,X_{d})$ is the orthogonal projection of $(X_{1},\ldots,X_{d})$ onto $\mathbb{V}_{k}$ when embedded in $\mathbb{V}_{n}$. Thus, $(V^{\top}X_{1}V,\ldots,V^{\top}X_{d}V)=\mathcal{P}_{k}(U^{\top}X_{1}U,\ldots,U^{\top}X_{d}U)\in\mathcal{C}_{k}$ since $\mathcal{C}_{n}$ is $\mathsf{O}_{n}$-invariant and since $\{\mathcal{C}_{n}\}$ is projection-compatible. ∎
@@ 4.5 Spectral Functions, (Quantum) Entropy, and Variants
Here $\mathrm{dom}(h_{n})=\Delta^{n-1}$ and $\mathrm{dom}(D_{n})=(\mathbb{R}^{n}_{+})^{2}$, while $\mathrm{dom}(\mathsf{H}_{n})=\mathcal{D}^{n-1}$ is the spectraplex from Example 1.1(b) and $\mathrm{dom}(S_{n})=(\mathbb{S}^{n}_{+})^{2}$. We use the standard convention that $0\log\frac{0}{y}=0$ even if $y=0$, and $x\log\frac{x}{0}=\infty$ when $x\neq 0$ [72, §2.3]. These sequences of functions are intersection-compatible but not projection-compatible (e.g., their domains are not projection-compatible).
@@
$E_{n}^{(m,k)}$ $=\Bigg\{(X,t)\in\mathbb{S}^{n}\oplus\mathbb{R}\Bigg|\exists T_{0},\ldots,T_{m},Z_{0},\ldots,Z_{k}\in\mathbb{S}^{n}\textrm{ s.t. }Z_{0}=I_{n},\ \sum_{j=1}^{m}w_{j}T_{j}=-2^{-k}T_{0},$
$\qquad\begin{bmatrix}Z_{i}&Z_{i+1}\\ Z_{i+1}&X\end{bmatrix}\succeq 0,\textrm{ for }i=0,\ldots,k-1,\ \begin{bmatrix}Z_{k}-X-T_{j}&-\sqrt{s_{j}}T_{j}\\ -\sqrt{s_{j}}T_{j}&X-s_{j}T_{j}\end{bmatrix}\succeq 0,$
$\qquad\textrm{for }j=1,\ldots,m,\ \mathrm{Tr}(T_{0})\leq t\Bigg\},$
@@
This is a dimension-free description of the form (ConicSeq) which almost, but not quite, satisfies the conditions of Proposition 3.2. Indeed, let $\mathscr{W}=\mathscr{V}^{\oplus(m+k+1)}$ and $\mathscr{U}=\mathscr{V}\oplus(\mathbb{S}^{2}\otimes\mathscr{V})^{\oplus(m+k)}\oplus\mathbb{R}$ containing the cones $\{\mathcal{K}_{n}=\{0\}\oplus(\mathbb{S}^{2}\otimes\mathbb{S}^{n})_{+}^{\oplus(m+k)}\oplus\mathbb{R}_{+}\}$. Define
@@
$\hskip 18.49988pt\bigoplus_{j=1}^{m}\left(\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes Z_{k}-\begin{bmatrix}1&\sqrt{s_{j}}\\ \sqrt{s_{j}}&s_{j}\end{bmatrix}\otimes T_{j}\right),-\mathrm{Tr}(T_{0})\Bigg),$
@@
Note that $\{A_{n}\}$ and $\{B_{n}\}$ are morphisms, that $\{u_{n}\}$ is a freely-described element of $\mathscr{U}$ satisfying $u_{n+1}-u_{n}\in\mathcal{K}_{n+1}$, but that $\{B_{n}^{*}\}$ is not a morphism because of the $\mathrm{Tr}(T_{0})$ term in $B_{n}$. By the proof of Proposition 3.2, this description certifies that $E_{n}^{(m,k)}\subseteq E_{n+1}^{(m,k)}$ but not $E_{n+1}^{(m,k)}\cap(\mathbb{S}^{n}\oplus\mathbb{R})\subseteq E_{n}^{(m,k)}$. Analogously to the naïve description of the solid simplex in Section 4.1, this dimension-free description does not make the intersection compatibility of $\{E_{n}^{(m,k)}\}$ obvious.
@@
Note that $\mathbb{S}^{2}\otimes\mathbb{S}^{n}\cong(\mathbb{S}^{n})^{3}$ as $\mathsf{O}_{n}$-representations, and these isomorphisms commute with zero-padding, so that $\mathscr{U}\cong\mathscr{V}^{1+3(m+k)}\oplus\mathbb{R}$ as consistent sequences. As $\dim(\mathbb{S}^{n})^{\mathsf{O}_{n}}=1$ and $\dim\mathcal{L}(\mathbb{S}^{n})^{\mathsf{O}_{n}}=2$, the dimension of invariants parametrizing dimension-free descriptions are
$\dim\mathcal{L}(\mathbb{V}_{n},\mathbb{U}_{n})^{\mathsf{O}_{n}}=3[2(m+k)+1],\qquad\dim\mathcal{L}(\mathbb{W}_{n},\mathbb{U}_{n})^{\mathsf{O}_{n}}=3(m+k+1)[2(m+k)+1],$
@@ 4.6 Graph Parameters
Let $\mathscr{V}=\{\mathbb{S}^{n}\}$ with embeddings by zero-padding and the action of $\mathsf{G}_{n}=\mathsf{S}_{n}$ by conjugation. As discussed in Section 1.1.2 and Example 1.7, graph parameters are sequences $\{f_{n}\colon\mathbb{S}^{n}\to\mathbb{R}\}$ of $\mathsf{G}_{n}$-invariant functions, and many standard graph parameters are convex (or concave) and satisfy either intersection or projection compatibility. Furthermore, it is desirable to find relaxations for standard graph parameters satisfying the same compatibility conditions. We consider here the examples of max-cut and inverse stability number, and we show that their usual relaxations are freely-described with descriptions certifying their compatibility. We then derive parametric families of dimension-free descriptions from our framework that may be used to fit graph parameters to data.
@@
The sequence $\{\mathcal{E}_{n}\}$ also satisfies projection compatibility, and the above is a dimension-free description certifying this compatibility. Indeed, let $\mathscr{W}=\{0\}$ and $\mathscr{U}=\mathscr{V}\oplus\{\mathbb{R}^{n}\}$ where the latter sequence is the usual one from Example 1.5, with cones $\mathscr{K}=\{\mathbb{S}^{n}_{+}\oplus\{0\}\}$. Then (17) is of the form (ConicSeq) with $A_{n}X=(X,\mathrm{diag}(X))$, $B_{n}=0$, and $u_{n}=(0,-\mathbbm{1}_{n})$. Note that $\{A_{n}\},\{A_{n}^{*}\}$ are morphisms and $u_{n+1}-u_{n}=(0,-e_{n+1})=(e_{n+1}e_{n+1}^{\top},0)-A_{n+1}e_{n+1}e_{n+1}^{\top}$, hence Proposition 3.2(b) applies. Neither the cut polytopes nor the elliptopes is intersection-compatible, as zero-padding a matrix with all-1’s diagonal does not yield such a matrix. The sequences $\{\mathrm{CUT}_{n}-I_{n}\}$ and $\{\mathcal{E}_{n}-I_{n}\}$ are, however, both intersection- and projection-compatible, and the shifted elliptopes admit dimension-free descriptions certifying their compatibility.
@@
where $X\geq 0$ denotes an entrywise nonnegative matrix. Both $\{\mathcal{D}_{n}\}$ and $\{\widetilde{\mathcal{D}}_{n}\}$ are intersection-compatible. Moreover, the above description of $\{\widetilde{\mathcal{D}}_{n}\}$ is dimension-free and certifies this compatibility as in Proposition 3.2(b). Indeed, let $\mathscr{W}=\{0\}$ and $\mathscr{U}=\mathscr{V}\oplus\mathbb{R}$ with cones $\mathcal{K}_{n}=(\mathbb{S}^{n}_{+}\cap\mathbb{R}^{n\times n}_{+})\oplus 0$. Then the above description of $\widetilde{\mathcal{D}}_{n}$ is of the form (ConicSeq) with $A_{n}X=(X,\mathbbm{1}_{n}^{\top}X\mathbbm{1}_{n})$, $B_{n}=0$, and $u_{n}=(0,-1)$. Note that $\{A_{n}\}$ is a morphism and $u_{n}=u_{n+1}$, hence Proposition 3.2(a) applies. Neither $\mathcal{D}_{n}$ nor $\widetilde{\mathcal{D}}_{n}$ is projection-compatible since their support functions, which are the inverse stability number and its semidefinite approximation above, are not intersection compatible (see Section 1.3). Indeed, appending isolated vertices to a graph increases its stability number.
@@
Beyond the two particular preceding examples, we derive here an expressive parametric family of convex graph parameters; in addition to above examples, our family also includes several other examples from [74]. To that end, let $\mathscr{W}=\mathscr{U}=\mathrm{Sym}^{2}(\mathrm{Sym}^{\leq 2}\{\mathbb{R}^{n}\})=\{\mathbb{S}^{\binom{n+2}{2}}\}$. The dimensions of invariants parametrizing dimension-free descriptions in this case is too large for us to write explicit bases for them as functions of $n$. Instead, we compute these dimensions using the algorithm in Section 6.1 below, see Example 6.2(b):
@@
[…]